A freight train and an express train leave towns 390km apart, traveling toward one another. The freight train travels 30km per hr slower than the express train. They pass one another 3hr later. What are their speeds?
Answer by checkley77(9402)

390=3x+3(x+30)
390=3x+3x+90
6x=390-90
6x=300
x=300/6
x=50 kmh. is the speed of the freight train.
50+30=80 kmh. for the speed of the express train.
Proof:
390=3*50+3*80
390=150+240
390=390